Tanning Company analyzes its receivables to estimate bad debt expense. The accounts receivable balance is $300,000 and credit sales are $1,300,000. An aging of accounts receivable shows that approximately 5% of the outstanding receivables will be uncollectible. What adjusting entry will Tanning Company make if Allowance for Doubtful Accounts has a credit balance of $2,500 before adjustment? Bad Debt Expense = 12,500 Allowance for Doubtful Accounts = 12,500 Bad Debt Expense = 19,500 Allowance for Doubtful Accounts = 19,500 Bad Debt Expense = 22,000 Allowance for Doubtful Accounts = 22,000 Bad Debt Expense = 65,000 Allowance for Doubtful Accounts = 65,00.
Answer #1
Estimated uncollectible = 300000*5% = $15000
Bad Debt Expense = 15000-2500 =$12500
The adjusting entry fo Bad Debts expense is:
Bad Debt Expense 12500
      Allowance for Doubtful Accounts 12500